What is the locus of P if it`s polar with respect to the hyperbola (x^2/a^2-y^2/a^2)=1 be equally inclined to the coordinate axes?

Grade:12th pass

1 Answers

Deepak Kumar Shringi
6 years ago
equation of a polar is always T=0 with respect to the pole here polar is equally inclined to axes that means slope is -1,1
now T=0
xx1/a2 – yy1/b2=1
slope is x1b2/y1a2=1 or -1
then x1/y1= a2/b2